# Federated Learning Privacy: Attacks, Defenses, Applications, and Policy Landscape — A Survey

### v1

@@ 2.5. Metrics measuring privacy
Membership inference attacks - accuracy. The goal of membership inference attacks is to infer whether a particular instance is used in the training of a target model. A balanced evaluation set, which consists of the same number of members and non-members, is used in the membership inference attack evaluations in most existing literature. For membership inference attacks, the attack accuracy measures the attack effectiveness over the whole data population (Shokri et al., 2017). Attack AUC can also be used to measure the privacy leakage in terms of membership inference over the whole population. The higher the attack accuracy is, the more privacy leakage the target model has on its training set for the whole set.
@@
Membership inference attacks - TPR at a low FPR. Accuracy and AUC measures the average privacy vulnerability across all instances. In the context of privacy, one usually wants to ensure that every individual’s privacy is protected. It is thus necessary to measure membership leakage threat for the most vulnerable instances. In (Carlini et al., 2022), it was proposed to use true positive rate (TPR) at a low false positive rate (FPR) to measure effects of MI attacks. Consider, for example, TPR at FPR. If is also around , then that means the adversary cannot do better than random guessing even when they are very confident about MI inference. On the other hand, if If is also around , then that means the adversary would be able to correctly identify (a small percentage of) members with a odds ratio.
@@
Data reconstruction attacks - reconstruction quality. One important metric for gauging the success of data reconstruction attacks is in the similarity of the reconstructions to the ground truth images. This measure can be visual and is typically quantified through image similarity metrics such as PSNR (peak signal-to-noise ratio), SSIM (structural similarity index measure), or L-PIPS (learned perceptual image patch similarity). Image identifiability precision (IIP) was also introduced as a potential metric for reconstruction quality (Yin et al., 2021a). Instead of per-image calculatioins, compared to all images in the ground truth batch, IIP finds the nearest neighbor of each reconstructed image. The metric is quantified as the fraction of reconstructions that have the ground truth image as their nearest neighbor. Additionally, (Zhao et al., 2024) discusses the use of leaked data for training models in downstream prediction tasks related to the image similarity metrics. Even reconstructed images with some of the lowest metric scores (e.g., PSNR, SSIM) were still shown to improve the accuracy of the trained models (Zhao et al., 2024).
@@
Data reconstruction attacks - leakage rate. Another metric used to quantify success of data reconstruction attacks is leakage rate. While leakage rate can be simply defined as the number of leaked images, quantification can still vary. Some methods such as linear layer leakage can result in leaking a single image with multiple ground truth images overlapping. Here, leakage rate is typically defined as the number of reconstructions that only have a single image leaked. While methods such as gradient inversion typically rely on image similarity to measure the quality, leakage rate can also be used based on the number of images above a given threshold for any metric (e.g., PSNR) (Hatamizadeh et al., 2023).
@@ 3. Relation to Prior Surveys
Federated learning was originally introduced as a method to collaboratively train machine learning models while allowing users to keep their data local and private. Despite this, many prior works have shown that the updates sent to a server still inherently include sensitive information that can be used to leak either properties of the data or reconstruct them directly. Following this, much work has also been done in the area of FL privacy defense. To motivate our work, we first compare the prior surveys in FL privacy and discuss the missing pieces. Table 1 shows a summary of prior surveys compared to this work.
@@
Many surveys have looked at the privacy aspect of FL. (El Ouadrhiri and Abdelhadi, 2022; Khan et al., 2023; Li et al., 2020) discuss privacy defense in FL with little discussion on privacy attacks. Other surveys (Mothukuri et al., 2021; Zhang et al., 2022; Yin et al., 2021b; Bouacida and Mohapatra, 2021; Chen et al., 2022; Sikandar et al., 2023; Liu et al., 2022a) look at privacy attacks, but are missing discussion on more recent catastrophic attack methods such as linear layer leakage for data reconstruction or membership inference attacks that don’t require training data. (Chen et al., 2023b) discuss the trade-off between privacy and fairness in FL. While the methods of optimization and closed-form linear layer leakage attacks are discussed, the work still misses other key data reconstruction methods such as multi-round disaggregation or attacks targeted against secure aggregation. Similarly, (Rigaki and Garcia, 2023) discusses privacy attacks in machine learning but also lacks many current privacy attacks specific to FL. (Lyu et al., 2022) is the closest survey to ours, focusing on security and privacy attacks and defenses in FL. However, important discussion on the key limitations of attacks and defenses is missing. Similarly, discussion into recently introduced data reconstruction attacks such as (Fowl et al., 2022; Wen et al., 2022; Lam et al., 2021; Kariyappa et al., 2023) are also missing.
@@
With the rapid development of FL privacy attacks and defenses, prior surveys are missing discussion of many current key privacy attacks and defenses. Our work bridges this gap and provides a comprehensive discussion on the current privacy attack threats (data reconstruction, membership-inference, and property inference attacks) along with defenses (differential privacy, secure aggregation, homomorphic encryption, and trusted execution environments) while highlighting the limitations of these methods along with areas where improvement is needed. Additionally, to the best of our knowledge, we are the only work to discuss current industry applications utilizing FL in a variety of industry sectors including healthcare, finance, and IoT edge applications. Similarly, we are the only work to include crucial discussion on government data privacy policies and their relation to FL. We close our work by looking ahead and hypothesizing the key developments that would be necessary for FL to become a privacy-preserving technology that addresses user privacy concerns and policy mandates in real-world deployments.
@@ 4. Application Use Cases and Policy Drivers
As part of the research in FL, investigating when the participating parties are adequately incentivized to participate is crucial, especially for real-life scenarios. The incentive mechanism expects that not all clients have the same interests in the global model as the server. The existing FL incentive mechanism is typically designed by sampling a fixed subset of clients based on their data quantity or resources (Luo et al., 2023). On the far-end of this spectrum, on the non-federated setting, clients can train their own local model using only the available local data and send the local model parameters to the server. Analysis based on (Mansour et al., 2020) shows that local training often outperforms federated models, depending on the local data . Federated learning has multiple applications across various industries, including healthcare, finance, automotive, and personalized medicine. In healthcare, we can use federated learning to train models on patients’ data without compromising their privacy, especially when using Differential Privacy. Similarly, in the financial industry, we can employ federated learning to train fraud detection models without exposing any sensitive financial data to potentially malicious entities. In the automotive industry, we can employ federated learning to train autonomous driving models on real-life data avoiding transferring huge volumes of data back to the servers.
@@ 4.1. Application of FL in Healthcare
In the past few years, we have experienced a flood of changes in automation in the healthcare industry, increasing the need and expectation for better ML models. Healthcare professionals have turned to technology to help them treat patients better, improve their understanding of the emerging global demand, etc. However, training healthcare-related models requires vast and diverse datasets. On the other hand, sharing information is becoming even more challenging due to the ever-stricter privacy regulations. Using FL, participating institutions can train a global model on their in-house data without exposing sensitive local data. This new paradigm leads to better models and more accurate training algorithms, allowing them to access and also work around various regulations. Thus, the potential benefits of FL in the healthcare industry are significant.
@@
There have been several real-world efforts around FL in the healthcare sector, including proof-of-concept projects and establishing long-running consortia (Rieke et al., 2020; Rehman et al., 2023). For example, in the United Kingdom, FLIP enables AI researchers to develop clinical applications using NHS patient data without transferring the information outside the hospital network, addressing the challenge of confidentiality in healthcare (FLI, [n. d.]). In France, the HealthChain consortium (hea, [n. d.]) aims to address the bottleneck in accessing medical data for AI models that could lead to medical breakthroughs. In Germany, the Trustworthy Federated Data Analytics (TFDA) project (TFD, [n. d.]) aims to address the challenges posed by centralized data structures in the context of growing data needs for machine learning. Recognizing the drawbacks and threats associated with data centralization, TFDA focuses on implementing decentralized, cooperative data analytics architectures, specifically within and beyond the Helmholtz research community. The MELLODDY project (Heyndrickx et al., 2023; Oldenhof et al., 2023), co-funded by the European Union and EFPIA companies, has successfully demonstrated the feasibility of large-scale collaborative artificial intelligence (AI) for drug discovery. By leveraging a large collection of small molecules, the project utilized federated and privacy-preserving machine learning to achieve more accurate predictive modeling, leading to potential efficiencies in drug discovery. The German Cancer Consortium’s Joint Imaging Platform (JIP) (Scherer et al., 2020) is a strategic initiative within the German Cancer Consortium (DKTK) with the goal of establishing a technical infrastructure to facilitate modern and distributed imaging research. Emphasizing the use of contemporary machine learning methods in medical image processing, the project aims to enhance collaboration among clinical sites and support multicenter trials. Internationally, the Federated Tumor Segmentation (FeTS) initiative (fet, [n. d.]; Pati et al., 2021) is developing one of the largest federations of healthcare institutions. This initiative aims to gain knowledge for tumor boundary detection from diverse patient populations without sharing patient data.In a collaborative effort involving seven clinical institutions worldwide, an early study explored using FL to construct robust medical imaging classification models for breast density based on BI-RADS (Roth et al., 2020). Despite significant differences in datasets across sites, including mammography systems, class distribution, and data set size, the results demonstrated successful AI model training in a federated manner. The EXAM project (Dayan et al., 2021) aimed to build a common predictive model to estimate the oxygen treatment requirements of patients arriving at the emergency department (ED) with symptoms of COVID-19 for the next 24h and 72h periods. A consortium of 20 hospital sites across the globe participated in this real-world application of FL. Furthermore, institutions such as The American College of Radiology (ACR) enable privacy-preserving AI utilizing federated learning technology (acr, [n. d.]).
These efforts show the trend within the healthcare industries to decentralize data for AI workflows while keeping regulatory challenges in mind for the benefit of patient privacy. FL technology might be the key to enabling trustworthy AI development, such as diversity (Commission et al., 2019).
@@ 4.2. Application of FL in Finance
The data privacy and protection laws are becoming ever more restrictive, especially in industries like the Financial sector. This allows consumers and businesses to trust each other to keep the data safe and secure. With traditional ML, businesses dependent on FinTech face several issues, such as getting clearance and lawful consent, the governance of the customers’ data and the time and finally, the cost in collecting and of managing such data. FL provides an efficient way for managing this data by keeping it local, on the financial institutions servers. FL is an encrypted and distributed machine learning approach that allows joint training on decentralized data where participants do not share any of it.
@@
One of the most promising application of federated learning in financial services is anti-money laundering (AML). In a typical money laundering scheme, criminals try to conceal the origin, identity, and destination of financial transactions to look normal as they are from a legitimate source. AML solutions allow financial institutions to prevent, detect, investigate and report any activity that is suspicious of money laundering and stop criminals from illegally obtained funds as legitimate income. The estimated amount of money laundered globally in one year is 2 - 5% of global GDP ($800 billion - $2 trillion)(UN-, [n. d.]). Failure to prevent money laundering can endanger the integrity and stability of global financial system. Financial institutions are required to put strict policies and measures to prevent, detect, and report laundered money and combat these crimes. However, each financial institution has typically access to its own transnational data and can not prevent and detect complex money laundry cases where the money routes through multiple financial institutions that could look legitimate to some of involved institutions in the chain. Cross-silo Federated AML can greatly help financial institutions to create better predictive models with broader context for identifying bad actors. Such models learn more global and deeper correlation identifiers for bad actors and money laundry specific actions based on contributed data provided by consortium of financial institutions.
@@
Another application of federated learning is in insurance underwriting. The main use cases for federated learning in insurance include origination- predicting if a person contracts an insurance offer, insurer’s risk assessment, fraud detection, and claim processing. Due to soiled data in different companies on different demographics or customers movement across insurance companies, Insurers can significantly improve their predictive models for these use cases by utilizing federated learning. Claim history processing is of the most challenging use cases that federated learning can lead to more accurate and efficient operations. Insurance companies often need to have access to previous claims history of an insurer in details to give a better pricing or benefits to new clients. However, having access to such historical data in a privacy preserving manner is challenging due to regulations governing the insurance industry. Insurance companies can leverage cross-silo federated learning to privately train their models without compromising customers privacy and satisfying the governing regulations. Such models lead to more accurate offers while helping customers get the benefits of having a long-term history of free claims across different categories.

### v2

@@ 2.6. Metrics measuring privacy
Membership inference attacks - accuracy. The goal of membership inference attacks is to infer whether a particular instance is used in the training of a target model. A balanced evaluation set, which consists of the same number of members and non-members, is used by the membership inference attack evaluations in most existing literature. For membership inference attacks, the attack accuracy measures the attack effectiveness over the whole data population (Shokri et al., 2017). Attack AUC can also be used to measure the privacy leakage in terms of membership inference over the whole population. The higher the attack accuracy is, the more privacy leakage the target model has on its training set for the whole set.
@@
Membership inference attacks - TPR at a low FPR. Accuracy and AUC measures the average privacy vulnerability across all instances. In the context of privacy, one usually wants to ensure that every individual’s privacy is protected. It is thus necessary to measure membership leakage threat for the most vulnerable instances. In (Carlini et al., 2022), it was proposed to use true positive rate (TPR) at a low false positive rate (FPR) to measure effects of MI attacks. Consider, for example, TPR at FPR. If is also around , then that means the adversary cannot do better than random guessing even when they are very confident about MI inference. On the other hand, if If is around , then an adversary would be able to correctly identify (a small percentage of) members with a odds ratio.
@@
Data reconstruction attacks - reconstruction quality. One important metric for gauging the success of data reconstruction attacks is in the similarity of the reconstructions to the ground truth images. This measure can be visual and is typically quantified through image similarity metrics such as PSNR (peak signal-to-noise ratio), SSIM (structural similarity index measure), or L-PIPS (learned perceptual image patch similarity). Image identifiability precision (IIP) was also introduced as a potential metric for reconstruction quality (Yin et al., 2021a). Instead of per-image calculatioins, compared to all images in the ground truth batch, IIP finds the nearest neighbor of each reconstructed image. The metric is quantified as the fraction of reconstructions that have the ground truth image as their nearest neighbor. Additionally, (Zhao et al., 2024) discusses the use of leaked data for training models in downstream prediction tasks related to the image similarity metrics, where even reconstructed images with the lowest metric scores (e.g., PSNR, SSIM) were still shown to improve the accuracy of trained models (Zhao et al., 2024).
@@
Data reconstruction attacks - leakage rate. Another metric used to quantify success of data reconstruction attacks is leakage rate. While leakage rate can be simply defined as the number of leaked images, quantification can still vary. Some methods, such as linear layer leakage, can result in leaking an image consisting of multiple client images overlapping. Here, leakage rate is typically defined as the number of reconstructions with only have a single, non-overlapping image leaked. Although methods such as gradient inversion typically rely on image similarity to measure quality, leakage rate can also be used based on the number of images above a given threshold for any metric (e.g., PSNR) (Hatamizadeh et al., 2023).
@@ 3. Relation to Prior Surveys
Federated learning was originally introduced as a method to collaboratively train machine learning models while allowing users to keep their data local and private. Despite this, many prior works have shown that the updates sent to a server still inherently include sensitive information that can be used to leak either properties of the data or reconstruct them directly. Following this, much work has also been done in the area of FL privacy defense. To motivate our work, we first compare the prior surveys in FL privacy and discuss the missing pieces. Table 1 shows a summary of previous surveys compared to this work.
@@
Many surveys have looked at the privacy aspect of FL. The works of (El Ouadrhiri and Abdelhadi, 2022; Khan et al., 2023; Li et al., 2020) discuss privacy defense in FL with little discussion on privacy attacks. Other surveys (Mothukuri et al., 2021; Zhang et al., 2022; Yin et al., 2021b; Bouacida and Mohapatra, 2021; Chen et al., 2022; Sikandar et al., 2023; Liu et al., 2022a) look at privacy attacks, but are missing discussion on more recent catastrophic attack methods such as linear layer leakage for data reconstruction or membership inference attacks that do not require training data. Chen et al. (2023b) discusses the trade-off between privacy and fairness in FL. While the methods of optimization and closed-form linear layer leakage attacks are discussed, the work still misses other key data reconstruction methods such as multi-round disaggregation or attacks targeted against secure aggregation. Similarly, Rigaki and Garcia (2023) discusses privacy attacks in machine learning but also lacks many current privacy attacks specific to FL. Lyu et al. (2022) is the closest survey to ours, focusing on security and privacy attacks and defenses in FL. However, important discussion on the key limitations of attacks and defenses is missing. Similarly, recent, powerful data reconstruction attacks such as (Fowl et al., 2022; Wen et al., 2022; Lam et al., 2021; Kariyappa et al., 2023) are missing.
@@
Our work bridges this gap on the recent and rapid development of privacy attacks and defenses and provides a comprehensive discussion on current privacy threats (data reconstruction, membership-inference, and property inference attacks) and defenses (differential privacy, secure aggregation, homomorphic encryption, and trusted execution environments) while highlighting the limitations of these methods along with areas where improvement is needed. Additionally, to the best of our knowledge, we are the only work that discusses current industry applications utilizing FL in a variety of industry sectors including healthcare, finance, and IoT edge applications. Similarly, we are the only work to include crucial discussion on government data privacy policies and their relation to FL. We close our work by looking ahead and hypothesizing the key developments that would be necessary for FL to become a privacy-preserving technology that addresses user privacy concerns and policy mandates in real-world deployments.
@@ 4. Application Use Cases and Policy Drivers
Federated learning has multiple applications across various industries, including healthcare, finance, automotive, and IoT/Edge devices. In healthcare, FL can be used to train models on patients’ data without compromising privacy, especially when using differential privacy. Similarly, in the financial industry, we can employ FL to train fraud detection models without exposing sensitive financial data to potentially malicious entities. In the automotive industry, FL can be employed to train autonomous driving models on real-life data while avoiding the costs of transferring huge volumes of data back to the servers. Table 2 further summarizes the application domains along with the commonly applicable setting (cross-device and cross-silo) for deployment.
FL has been successfully deployed in industry settings (Ge et al., 2021) (Imp, 2021) (Zhang et al., 2021a) (Sun et al., 2024). One example of successful adoption comes from predictive maintenance and failure prediction, which are crucial for large-scale production lines. They require real-time failure monitoring and updates from multiple resources across an enterprise that are often guarded by different privacy and security policies. Bosch (Ge et al., 2021) has used many federated algorithms in various horizontal FL and vertical FL production scenarios such as predictive maintenance. Through this, Bosch was able to improve the life expectancy of production line machinery over time without significant overhead compared to centralized learning approaches. Other companies (Imp, 2021) have also used FL in similar use cases to drastically reduce the cost of halting in the production line due to unexpected failures.
@@ 4.1. Application of FL in Healthcare
Over the past few years, the automation of the healthcare industry has surged, driving an increasing demand for advanced ML models. Healthcare professionals are leveraging technology to enhance patient care and better meet emerging global demands while managing vast and sensitive datasets. Training effective healthcare models, however, requires access to diverse data sources, and sharing such information is increasingly challenging under strict privacy regulations. FL offers a promising solution by enabling collaborative model training across institutions without exposing sensitive local data. With FL, participating organizations can develop a global model based on their in-house data, leading to more accurate and diverse ML models (Commission et al., 2019) while complying with regulatory frameworks and mitigating data silo issues.
Several real-world efforts have been made around FL in the healthcare sector, including proof-of-concept projects and the establishment of long-running consortia (Rieke et al., 2020; Rehman et al., 2023). For example, in the United Kingdom, the FLIP project empowers AI researchers to develop clinical applications using NHS patient data without transferring the information outside the hospital network, thus addressing the challenge of confidentiality in healthcare (FLI, [n. d.]). In France, the HealthChain consortium (hea, [n. d.]) aims to address the bottleneck in accessing medical data for AI models that could lead to medical breakthroughs. In Germany, the Trustworthy Federated Data Analytics (TFDA) project (TFD, [n. d.]) aims to address the challenges posed by centralized data structures in the context of growing need for more data for machine learning. Recognizing the drawbacks and threats associated with data centralization, TFDA focuses on implementing decentralized cooperative data analytics architectures, specifically within and beyond the Helmholtz research community.
@@
The MELLODDY project (Heyndrickx et al., 2023; Oldenhof et al., 2023), co-funded by the European Union and EFPIA companies, has successfully demonstrated the feasibility of large-scale collaborative artificial intelligence (AI) for drug discovery. By leveraging a large collection of small molecules, the project utilized federated and privacy-preserving machine learning to achieve more accurate predictive modeling, leading to potential efficiencies in drug discovery. The German Cancer Consortium’s Joint Imaging Platform (JIP) (Scherer et al., 2020) is a strategic initiative within the German Cancer Consortium (DKTK) with the goal of establishing a technical infrastructure to facilitate modern and distributed imaging research. Emphasizing the use of contemporary machine learning methods in medical image processing, the project aims to enhance collaboration among clinical sites and support multicenter trials. Internationally, the Federated Tumor Segmentation (FeTS) initiative (fet, [n. d.]; Pati et al., 2021) is developing one of the largest federations of healthcare institutions. This initiative aims to gain knowledge for tumor boundary detection from diverse patient populations without sharing patient data.
In a collaborative effort involving seven clinical institutions worldwide, an early study explored using FL to construct robust medical imaging classification models for breast density based on BI-RADS (Roth et al., 2020). Despite significant differences in datasets across sites, including mammography systems, class distribution, and data set size, the results demonstrated successful AI model training in a federated manner. The EXAM project (Dayan et al., 2021) aimed to build a common predictive model to estimate the oxygen treatment requirements of patients arriving at the emergency department (ED) with symptoms of COVID-19 for the next 24h and 72h periods. A consortium of 20 hospital sites across the globe participated in this real-world application of FL. Furthermore, institutions such as The American College of Radiology (ACR) enable privacy-preserving AI utilizing FL technology (acr, [n. d.]).
The concept of federation has deep roots in computational genomics (Chaterji et al., 2019). Early work on federating computational resources—such as CPUs, GPUs, and FPGAs—and datastores demonstrated that distributed processing within unified portals could simplify data management and minimize extensive data transfers. This pioneering research remains relevant today by enabling the efficient handling of large-scale genomic data while ensuring privacy and security, paving the way for advanced federated approaches in healthcare. In particular, the seminal work on federated compute infrastructures for genomics highlights challenges such as heterogeneous data formats and disparate computational environments. By standardizing data exchange protocols and developing interoperable interfaces, early federated systems achieved seamless integration of diverse resources and minimized the need for large-scale data transfers. These innovations not only enhanced data security through localized processing and robust access controls but also laid the groundwork for overcoming technical and regulatory barriers—a legacy that informs and strengthens current FL applications in healthcare. Furthermore, computer systems research toward optimizing databases such as SOPHIA (Mahgoub et al., 2019) has leveraged these federated principles to tackle dynamic workload challenges in real-world settings. The SOPHIA framework optimizes the reconfiguration of clustered NoSQL databases for applications like MG-RAST—a global-scale metagenomics repository—thereby ensuring high throughput and data availability despite fluctuating workloads. This integration of federated techniques with adaptive database management demonstrates how cross-silo federation can enhance scalability, performance, and resilience in modern database systems, making it a critical enabler for both computational genomics and healthcare applications.
An example of cross-device federated learning in the healthcare setting is action recognition performed on heterogeneous embedded devices (Jain et al., 2021). In healthcare, such methods can be applied to patient monitoring tasks, including fall detection or tracking physical activity during rehabilitation. Each device processes its local sensor or video data and contributes to a collaboratively trained global model without sharing raw data. This approach effectively addresses the challenges posed by devices with varying computational capabilities. In particular, asynchronous federated optimization allows each device to send its model updates to the server as soon as its local training is complete, rather than waiting for a synchronized round. This reduces idle time and mitigates the straggler problem—where slower devices hold back the overall training progress—thus accelerating the training process while maintaining robust model convergence.
@@ 4.2. Application of FL in Finance
The data privacy and protection laws are becoming ever more restrictive, especially in industries like the Financial sector. This allows consumers and businesses to trust each other to keep the data safe and secure. With traditional ML, businesses dependent on FinTech face several issues, such as getting clearance and lawful consent, the governance of the customers’ data, and finally, the cost in collecting and of managing such data. FL provides an efficient way for managing this data by keeping it local, on the financial institutions servers. FL is an encrypted and distributed machine learning approach that allows joint training on decentralized data where participants do not need to share it.
@@
One of the most promising applications in financial services is anti-money laundering (AML). In a typical money laundering scheme, criminals try to conceal the origin, identity, and destination of financial transactions to look normal as they are from a legitimate source. AML solutions allow financial institutions to prevent, detect, investigate and report any activity that is suspicious of money laundering and stop criminals from illegally obtained funds as legitimate income. The estimated amount of money laundered globally in one year is 2 - 5% of global GDP ($800 billion - $2 trillion) (UN-, [n. d.]). Failure to prevent money laundering can endanger the integrity and stability of global financial system. Financial institutions are required to put strict policies and measures to prevent, detect, and report laundered money and combat these crimes. However, each financial institution typically only has access to its own transnational data and can not prevent and detect complex money laundry cases where money routed through multiple financial institutions could look legitimate to individually involved institutions in the chain. Cross-silo Federated AML can greatly help financial institutions to create better predictive models with broader context for identifying bad actors. Such models learn more global and deeper correlation identifiers for bad actors and money laundering specific actions based on contributed data provided by a consortium of financial institutions.
@@
Another application of FL is in insurance underwriting. The main use cases of FL in insurance include origination- predicting if a person contracts an insurance offer, insurer’s risk assessment, fraud detection, and claim processing. Due to soiled data in different companies on different demographics or customers movement across insurance companies, insurers can significantly improve their predictive models for these use cases by utilizing FL. Claim history processing is of the most challenging use cases that can lead to more accurate and efficient operations. Insurance companies often need to have access to previous claims history of an insurer in details to give a better pricing or benefits to new clients. However, having access to such historical data in a privacy preserving manner is challenging due to regulations governing the insurance industry. Insurance companies can leverage cross-silo FL to privately train their models without compromising customers privacy and satisfying the governing regulations. Such models lead to more accurate offers while helping customers get the benefits of having a long-term history of free claims across different categories.